OpenStudy (anonymous):

What is the square root of -25?

OpenStudy (anonymous):

I need the imaginary number

OpenStudy (wil476003):

-[5] maybe

OpenStudy (anonymous):

i5

OpenStudy (anonymous):

that is the square root since i=square root -1

OpenStudy (anonymous):

Okay so it would be -5, but then the i makes it positive?

OpenStudy (welshfella):

sqrt 25 = 5, sqrt -1 = i

OpenStudy (anonymous):

no its i5

OpenStudy (wil476003):

u are right

OpenStudy (welshfella):

its usually written 5i

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

it doesn't matter either way

OpenStudy (anonymous):

I want to give you all medals :(

OpenStudy (phi):

i is short for \( \sqrt{-1} \) and the idea is \[ \sqrt{-1} \sqrt{-1} = -1 \] notice 5i means 5*i and \[ (5i)^2 = 5 \cdot i \cdot 5 \cdot i = 5 \cdot 5 \cdot \sqrt{-1}\cdot \sqrt{-1} \\=25\cdot -1 = -25 \]

OpenStudy (anonymous):

Okay, that makes sence to me

OpenStudy (anonymous):

but I also have \[3\sqrt{-49}\]

OpenStudy (phi):

people sometimes think -5 is the square root of -25 but -5*-5 = + 25 so we need to use i for square roots, always break it into -1 and then the positive number so -49 = -1 * 7*7 now you can do the square root of that

OpenStudy (anonymous):

And thats also were I get stuck, I have the others once it was just those too

OpenStudy (phi):

\[ \sqrt{-1 \cdot 7 \cdot 7}= \sqrt{-1} \cdot \sqrt{7\cdot 7} \]

OpenStudy (phi):

I assume you know how to "take out a pair" from the square root? for example, \( \sqrt{7\cdot7}= 7 \)

OpenStudy (anonymous):

Yes, I think so

OpenStudy (phi):

and any time you see \( \sqrt{-1}\) remember to use its "nickname" i

OpenStudy (anonymous):

okay, I never put realized it was the same, it makes so much seance now

OpenStudy (anonymous):

sense

OpenStudy (phi):

what do you get for \[ 3 \sqrt{-49} \]? remember the 3 "out front' means 3 times sqr(-49)

OpenStudy (anonymous):

Right so 16.3?

OpenStudy (phi):

how did you get that ?

OpenStudy (anonymous):

division?

OpenStudy (phi):

let's just look at \[ \sqrt{-49}\] do you now we can write -49 as -1 * 7 * 7 ?

OpenStudy (anonymous):

it doesnt work

OpenStudy (anonymous):

Wait....

OpenStudy (anonymous):

yes it does lol

OpenStudy (phi):

the idea of a square root is you can write \[ \sqrt{7 \cdot 7}= 7 \]

OpenStudy (anonymous):

do I need to simplify that?

OpenStudy (phi):

so with -49 write it as -1*7*7 now do the square root \[ \sqrt{-49}= \sqrt{-1 \cdot 7 \cdot 7} \] remember we can write this as \[ \sqrt{-1 \cdot 7 \cdot 7} = \sqrt{-1} \cdot \sqrt{ 7 \cdot 7}\]

OpenStudy (phi):

you should be able to simplify this now: \[ \sqrt{-1} \cdot \sqrt{ 7 \cdot 7}\]

OpenStudy (phi):

what letter should you use to replace the \( \sqrt{-1} \) ? and how do you simplify \( \sqrt{7 \cdot 7} \) ? (see up above)

OpenStudy (anonymous):

So it would be just 49?

OpenStudy (phi):

7*7 is 49 \( \sqrt{7\cdot 7} \text{ is } 7 \)

OpenStudy (phi):

Or if you are confused, you could use a calculator to find sqr(49) though it would be better if you can do this problem in your head: 7*7 is 49 and sqr(7*7) is7

OpenStudy (anonymous):

yup, thanks

OpenStudy (anonymous):

i gtg but thank you!!!!

OpenStudy (idealist10):

The answer is 5i.